# Gender Bias and Property Taxes

### v1

@@ Abstract
Gender bias distorts the economic behavior and outcomes of women and households. We investigate gender biases in property taxes. We analyze records of more than 100,000 property tax appeal hearings and more than 2.7 years of associated audio recordings, considering how panelist and appellant genders associate with hearing outcomes. We first observe that female appellants fare systematically worse than male appellants in their hearings. Second, we show that, whereas male appellants’ hearing outcomes do not vary meaningfully with the gender composition of the panel they face, those of female appellants’ do, such that female appellants obtain systematically lesser (greater) reductions to their home values when facing female (male) panelists. Employing a multi-modal large language model (M-LLM), we next construct measures of participant behavior and tone from hearing audio recordings. We observe markedly different behaviors between male and female appellants and, in the case of male appellants, we find that these differences also depend on the gender of the panelists they face (e.g., male appellants appear to behave systematically more aggressively towards female panelists). In contrast, the behavior of female appellants remains relatively constant, regardless of their panel’s gender. Finally, we show that female appellants continue to fare worse in front of female panels, even when we condition upon the appelant’s in-hearing behavior and tone. Our results are thus consistent with the idea that gender biases are driven, at least in part, by unvoiced beliefs and perceptions on the part of ARB panelists. Our study documents the presence of gender biases in property appraisal appeal hearings and highlights the potential value of generative AI for analyzing large-scale, unstructured administrative data.
@@ 1 Introduction
We examine gender concordance effects in the economically important context of property taxation, specifically in the administrative process related to property appraisal appeals (Nathan, Perez-Truglia and Zentner, 2025). Property taxes represent the second-largest source of tax revenue in the United States, contributing an estimated \$547 billion in 2019, more than doubling revenue from corporate income tax (Giaccobasso et al., 2022). Property taxes are particularly important for sub-national governments, which rely on revenue from property taxes to fund essential services such as schools, police, and parks.
@@
We study data on property appraisal appeal hearings obtained from Harris County, Texas (the Houston area) and associated audio recordings. Exploiting exogenous variation in the composition of ARB hearing panels attributable to the fact that panelists are randomly assigned to hearings, we examine how appeal outcomes for female and male appellants may differ depending on the gender composition of their hearing panel. We focus on gender concordance between appellants and panelists, documenting several notable findings. We calculate panelist-specific savings adjudication rates and document significant heterogeneity, beyond what could be explained by chance. Further, we calculate estimates of panelist-specific differences in savings adjudication rates by appellant gender, reflective of anti-female (pro-male) bias, and find that female panelists are systematically less likely to adjudicate in favor of a female appellant as compared to a male appellant. Next, undertaking more formal regression analyses, we find consistent results; we show that female appellants are systematically less likely to be awarded reductions in their home value and that this is particularly true when their ARB panel is comprised of females. The effects we report are substantial; we estimate that female appellants are approximately 4.2 p.p. less likely to obtain any reduction to their home value when faced with an all-female panel as compared to an all-male panel. Further, we estimate that, when facing an all-female (versus all-male) panel, female appellants obtain an approximate 33% smaller reduction to their initial home appraisal, on average. Given the average home appraisal value reduction across all hearings is approximately \$20,200, the gender concordance effect we observe translates to a roughly \$6,700 difference in final appraised value and an approximate \$140 difference in annual property taxes owed (the property tax rate in Harris County is presently 2.13%).
Next, leveraging unstructured audio recordings associated with more than 80,000 of these hearings, i.e., more than 2.7 years of audio, we employ a multi-modal large language model (M-LMM), Gemini 1.5 Flash, to annotate appellant and panelist behaviors and tones in each hearing – a task that would be prohibitively costly employing human analysts. Our goal is to understand why the observed gender differentials might arise. Examining differences in the prevalence of behaviors and tones across hearings by appellant and panel chair gender, we first observe that appellants exhibit much more variation in their tone and behavior than panelists. This can be explained, to some extent, by the fact that panelists receive formal training and are aware that their behavior and statements may be subsequently reviewed in court should the appellant decide to appeal their decision further.
@@ 2 Related Work
Of course, beyond the broad literature dealing with gender concordance effects, our paper relates and contributes to the literature studying what motivates households to appeal their property taxes. In this literature, Nathan, Perez-Truglia and Zentner (2025) show that both expected tax savings and filing frictions play significant roles, Giaccobasso et al. (2022) show that households are less likely to appeal if they believe that the government services funded by those taxes will provide more significant personal benefit to them, and Nathan, Perez-Truglia and Zentner (2024) show that households are willing to pay more in taxes if they believe other households are contributing their fair share. Some papers in this literature study property tax appeals heterogeneity by race, gender, and partisanship (Nathan, Perez-Truglia and Zentner, 2025, 2022; Avenancio-León and Howard, 2022; Holz, Novgorodsky and Simon, 2023). While these studies focus on what motivates the initial decision to appeal, our focus is on biases that affect the process itself, at the ARB hearings stage. Of particular relevance to our study, Holz, Novgorodsky and Simon (2023) describes the results of an experiment that involved sending emails to county appraisers, posing as homeowners, and randomly varying the homeowner’s gender. Those authors find that female appraisers are less likely to answer emails from a female homeowner. Although Holz, Novgorodsky and Simon (2023) does not focus on appeal hearings, as we do, their result is initially suggestive of possible negative gender concordance effects in the property tax context. It should be noted, however, that appraisers play a rather marginal role in hearing outcomes in our setting; despite appraisers’ presence at the hearings, final decisions are ultimately rendered by the ARB panel.
@@ 3.2.1 Structured Data
Each protest hearing record includes information on the protested property value assessment, as well as the revised assessment (if any) that emerged as an outcome of the hearing, i.e., the adjusted property value. We consider the difference between these two measures. Specifically, we construct a binary indicator of $I(Saved)$ and a continuous, non-negative measure of the percentage saved, i.e., $PctSaved$, where the latter reflects the percentage decline from the original assessed value. The former value thus reflects savings that appellants receive at the extensive margin, while the latter reflects savings at both the extensive and intensive margin. In approximately 3,400 cases, the final assessed property value exceeds the original property value assessment. This is a rare occurrence, which likely reflects that pieces of information came to light during the hearing that was detrimental to the appellant’s case, leading the ARB to realize that the initial assessment was too low. We omit these hearings from our analyses.
Each protest hearing record also includes the names of the ARB panelists who participated in the hearing, the appraiser, and the individuals who attended the formal hearing as the appellant(s). We parse the first and last names of each party from the hearing record. We then impute the gender of each based on their parsed first name. [^8] This imputation is based on Social Security Administration (SSA) data (Tzioumis, 2018; Kaplan, 2021). It should be noted that multiple parties may attend the hearing representing the property in question. In such cases, we gender code the first name that is listed. We focus our analysis on the first appellant listed because, unlike the other names of ARB panelists and the appraiser, which are standardized and consistently reported – suggesting these values are populated from an IT system – all appellant names are reported together in a single field, and the formatting is highly inconsistent – presumably reflecting manual data entry. We elaborate on the implied measurement error later and the steps we take to address that error leveraging gender detected based on voices in our audio recordings. Gender inferences are made based on empirical probabilities derived from the SSA data. For example, if a particular first name is held by 100 people in the United States SSA records, and 98 of those people were identified as female, then we would infer a 98% probability that the individual in our sample is female. Our approach employs a majority threshold in assigning gender. However, our results are robust to using stricter thresholds, e.g., 80%, 90%, 95%, and indeed they grow stronger. In turn, we use this gender information to construct an indicator of appellant gender, as well as a count-based measure of the number of $FemalePanelists$ that were present.
@@ 3.2.2 Audio Data
Appellant Sounds Confident 0.896 0.305
@@
Appellant Presents to Evidence 0.589 0.492
Appellant Sounds Hostile 0.111 0.314
@@
Appellant Sounds Assertive 0.932 0.252
@@
Appellant Interrupts Panelist 0.335 0.472
Appellant Asks Panelist to Speak Louder 0.065 0.246
@@
Appellant Asks Panelist to Repeat 0.348 0.476
@@
Panelist Sounds Confident 0.999 0.024
Panelist Sounds Defensive 0.006 0.078
Panelist Sounds Hostile 0.000 0.015
Panelist Raises Voice 0.000 0.016
Panelist Asks Appellant to Speak Louder 0.019 0.136
Panelist Asks Appellant to Repeat 0.133 0.339
Panelist Interrupts Appellant 0.104 0.305
Panelist Disagrees with Appellant 0.327 0.469
@@ 4 Descriptive Evidence
Assuming the assignment of ARB panelists is a true experiment, we can perform a simple OLS regression, relating the AnySaved dummy to a binary indicator of whether the ARB panel chair is female, absent any controls, and the result will bear a plausibly causal interpretation. Doing so, we obtain the results in Table 4, where we observe evidence of a negative gender concordance effect for female appellants, yet no evidence for any such effect among male appellants.
@@
We depict the resulting estimates in Figure 1, ordering panelists along the horizontal axis from least to most likely to adjudicate savings. The average probability that savings are adjudicated spans a range of approximately 40 percentage points, from the most ’critical’ to the most ’favorable’ ARB panelist.
@@
Even if ARB panelists exhibited no systematic bias in their decisions, a normal distribution of savings adjudication rates across panelists might nonetheless be expected due to random variation in case characteristics. Accordingly, to explore whether the observed distribution truly reflects panelist-specific biases in decision-making (i.e., whether the distribution truly differs from what we might expect due to chance), we perform a permutation test (Good, 2013). We implement our test via a repeated shuffling of the hearing outcome variable (i.e., AnySaved) over hearings to recover artificial, simulated ARB panelist-specific estimates of savings adjudication rates under the null hypothesis that adjudication rates are, in fact, homogeneous across panelists. That is, by breaking the relationship between the outcome and assigned ARB panelist identities, we effectively simulate the null distribution of our panelist-specific test statistics absent any panelist-specific bias. We can then compare the observed distribution of test statistics against that null distribution to arrive at an empirical p-value. This approach effectively retains the structure of panelist case assignments while enabling us to test the null hypothesis that panelists have no systematic effect on outcomes.
@@
In Figure 2, we plot the true set of estimates in red, atop the simulated estimates from 1,000 randomly shuffled samples shown in gray. The simulated distributions show what would result from the objective cases – i.e., the distribution of savings if ARB panelists did not exhibit systematic bias. It is readily apparent that the true distribution has much fatter tails than any simulated distribution, such that the variance in panelists’ rates of savings adjudication is wider than can be explained by chance alone.
Figure 2: Permutation test evaluating panelist bias; variance of true estimates yields an empirical p-value of 0.
@@ 5 Main Regression Results
That said, it is important to also recognize that several administrative restrictions limit how property taxes can be adjusted. These restrictions influence the incentives for households to appeal an appraisal, as well as the potential savings households may obtain from an appeal. Further, these restrictions make it difficult to evaluate the temporal evolution of appeals and savings arising from them (Giaccobasso et al., 2022). Most notably, property tax increases are capped at 10% annually. Households already under the cap have less incentive to appeal since their efforts would not immediately reduce their property taxes.
We next explore our measure of the dollar amount saved. As noted earlier, these estimations employ the logarithm of the amount saved as the outcome. The coefficient estimates are thus interpretable as approximate percentage effects. Considering Table 6, we observe the same general pattern of effects as in Table 5. Whereas female homeowners’ case outcomes differ systematically depending on the gender of the ARB panelists, outcomes for male homeowners do not. Considering our estimations in column 1, which employs the same right-hand side model as in column 1 of Table 5, we see that female homeowners save approximately 12.2% less (more) when faced with a female (male) ARB panel chair.131313exp(-0.13) - 1 = -0.122 Further, we see that the amount saved declines with the addition of more female panelists, such that the difference in home value reduction obtained by female homeowners when facing an all-female versus all-male panel is approximately 32.8%.141414exp(-0.397) - 1 = -0.328 Incorporating account fixed effects, we observe consistent findings in the case of ARB chair gender, though statistical significance declines slightly.
@@ 6 Audio Analysis
We next leverage a multi-modal large language model (LLM), specifically Gemini 1.5 Flash, to process all audio recordings associated with nearly all owner-attended ARB hearings in our original sample. Ultimately, we work with a subset of 80,197 audio recordings, namely those we can reliably match to hearings from our original sample of structured hearing records. Some structured hearing records in our sample do have associated recordings. Further, in some cases, we have multiple hearings for the same property in the same tax year (e.g., if separate hearings were held to adjudicate different protest reasons), and we lack any means of correctly discerning which recording associates with which hearing. Employing Gemini, we construct measures that quantify appellant behavior during a hearing. Using this data, we hope to understand i) any systematic differences that female homeowners may exhibit depending on the gender composition of the panel they face, and ii) any systematic differences in the relationship between homeowners’ behavior and hearing outcomes, depending on ARB panelist gender.
As noted earlier, we consider a measure of hearing length (in seconds), extracted from the metadata of each audio recording. Further, we consider Gemini’s assessment of the fraction of time an appellant spent talking during a hearing, binary indicators of how the appellant sounds, e.g., confident, frustrated, nervous, defensive, assertive, etc., and binary indicators of appellant actions during the hearing, e.g., referencing evidence, raising their voice, interrupting a panelist, disagreeing with a panelist, saying thank you, asking questions, asking a panelist to speak louder, or asking a panelist to repeat themselves.

### v2

@@ Abstract
Gender bias distorts the economic behavior and outcomes of women and households. We investigate gender biases in property taxes. We analyze records of more than 100,000 property tax appeal hearings and more than 2.7 years of associated audio recordings, considering how panelist and appellant genders associate with hearing outcomes. We first observe that female appellants fare systematically worse than male appellants in their hearings. Second, we show that, whereas male appellants’ hearing outcomes do not vary meaningfully with the gender composition of the panel they face, those of female appellants’ do, such that female appellants obtain systematically lesser (greater) reductions to their home values when facing female (male) panelists. Employing a multi-modal large language model (M-LLM), we next construct measures of participant behavior and tone from hearing audio recordings. We observe markedly different behaviors between male and female appellants and, in the case of male appellants, we find that these differences also depend on the gender of the panelists they face (e.g., male appellants appear to behave systematically more aggressively towards female panelists). In contrast, the behavior of female appellants remains relatively constant, regardless of their panel’s gender. Finally, we show that female appellants continue to fare worse in front of female panels, even when we condition upon the appellant’s in-hearing behavior and tone. Our results are thus consistent with the idea that gender biases are driven, at least in part, by unvoiced perceptions among ARB panelists. Our study documents the presence of gender biases in property appraisal appeal hearings and highlights the potential value of generative AI for analyzing large-scale, unstructured administrative data.
@@ 1 Introduction
We examine gender concordance effects in the economically important context of property taxation, specifically in the administrative process related to property appraisal appeals (e.g., Avenancio-León and Howard, 2022; Nathan, Perez-Truglia and Zentner, 2025). Property taxes represent the second-largest source of tax revenue in the United States, contributing an estimated \$547 billion in 2019, more than doubling revenue from corporate income tax (Giaccobasso et al., 2025). Property taxes are particularly important for sub-national governments, which rely on revenue from property taxes to fund essential services such as schools, police, and parks.
@@
We study data on property appraisal appeal hearings obtained from Harris County, Texas (the Houston area) and associated audio recordings. Exploiting exogenous variation in the composition of ARB hearing panels attributable to the fact that panelists are randomly assigned to hearings, we examine how appeal outcomes for female and male appellants may differ depending on the gender composition of their hearing panel. We focus on gender concordance between appellants and panelists, documenting several notable findings. We calculate panelist-specific savings adjudication rates and document significant heterogeneity, beyond what could be explained by chance. Further, we calculate estimates of panelist-specific differences in savings adjudication rates by appellant gender, reflective of anti-female (pro-male) bias, and find that female panelists are systematically less likely to adjudicate in favor of a female appellant as compared to a male appellant. Next, undertaking more formal regression analyses, we find consistent results; we show that female appellants are systematically less likely to be awarded reductions in their home value and that this is particularly true when their ARB panel is comprised of females. The effects we report are substantial; we estimate that female appellants are approximately 4.2 p.p. less likely to obtain any reduction to their home value when faced with an all-female panel as compared to an all-male panel. Further, we estimate that, when facing an all-female (versus all-male) panel, female appellants obtain an approximate 33% smaller reduction to their initial home appraisal, on average. Given the average home appraisal value reduction across all hearings is approximately \$20,200, the gender concordance effect we observe translates to a roughly \$6,700 difference in final appraised value and an approximate \$140 difference in annual property taxes owed (the average home in Harris County presently pays a tax rate of 2.13%).
Next, leveraging unstructured audio recordings associated with more than 80,000 of these hearings, i.e., more than 2.7 years of audio, we employ a multi-modal large language model (M-LMM), Gemini 1.5 Flash (Team, 2024), to annotate appellant and panelist behaviors and tones in each hearing – a task that would be prohibitively costly employing human analysts. Our goal is to understand why the observed gender differentials might arise. Examining differences in the prevalence of behaviors and tones across hearings by appellant and panel chair gender, we first observe that appellants exhibit much more variation in their tone and behavior than panelists. This can be explained, to some extent, by the fact that panelists receive formal training and are aware that their behavior and statements may be subsequently reviewed in court should the appellant decide to appeal their decision further.
@@ 2 Related Work
Of course, beyond the broad literature dealing with gender concordance effects, our paper relates and contributes to the literature studying what motivates households to appeal their property taxes. In this literature, Nathan, Perez-Truglia and Zentner (2025) show that both expected tax savings and filing frictions play significant roles, Giaccobasso et al. (2025) show that households are less likely to appeal if they believe that the government services funded by those taxes will provide more significant personal benefit to them, and Nathan, Perez-Truglia and Zentner (2024) show that households are willing to pay more in taxes if they believe other households are contributing their fair share. Some papers in this literature study property tax appeals heterogeneity by race, gender, and partisanship (Avenancio-León and Howard, 2022; Nathan, Perez-Truglia and Zentner, 2022, 2025; Holz, Novgorodsky and Simon, 2023). [^4] For other studies on property taxes, see, for example, Cabral and Hoxby (2012) and Kresch et al. (2023) While these studies focus on what motivates the initial decision to appeal, our focus is on biases that affect the process itself, at the ARB hearings stage. Of particular relevance to our study, Holz, Novgorodsky and Simon (2023) describes the results of an experiment that involved sending emails to county appraisers, posing as homeowners, and randomly varying the homeowner’s gender. Those authors find that female appraisers are less likely to answer emails from a female homeowner. Although Holz, Novgorodsky and Simon (2023) does not focus on appeal hearings, as we do, their result is initially suggestive of possible negative gender concordance effects in the property tax context. It should be noted, however, that appraisers play a rather marginal role in hearing outcomes in our setting; despite appraisers’ presence at the hearings, final decisions are ultimately rendered by the ARB panel.
@@ 3.2.1 Structured Data
Each protest hearing record includes information on the protested property value assessment, as well as the revised assessment (if any) that emerged as an outcome of the hearing, i.e., the adjusted property value. We consider the difference between these two measures. Specifically, we construct a binary indicator of $I(Saved)$ and a continuous, non-negative measure of the percentage saved, i.e., $PctSaved$, where the latter reflects the percentage decline from the original assessed value. The former value thus reflects savings that appellants receive at the extensive margin, while the latter reflects savings at both the extensive and intensive margin. In approximately 243 cases, the final assessed property value exceeds the original property value assessment. This is a rare occurrence, which likely reflects that pieces of information came to light during the hearing that was detrimental to the appellant’s case, leading the ARB to realize that the initial assessment was too low. We omit these hearings from our analyses.
Each protest hearing record also includes the names of the ARB panelists who participated in the hearing, the appraiser, and the individuals who attended the formal hearing as the appellant(s). We parse the first and last names of each party from the hearing record. We then impute the gender of each based on their parsed first name. [^9] This imputation is based on Social Security Administration (SSA) data (Kaplan, 2021). It should be noted that multiple parties may attend the hearing representing the property in question. In such cases, we gender code the first name that is listed. We focus our analysis on the first appellant listed because, unlike the other names of ARB panelists and the appraiser, which are standardized and consistently reported – suggesting these values are populated from an IT system – all appellant names are reported together in a single field, and the formatting is highly inconsistent – presumably reflecting manual data entry. We elaborate on the implied measurement error later and the steps we take to address that error leveraging gender detected based on voices in our audio recordings. Gender inferences are made based on empirical probabilities derived from the SSA data. For example, if a particular first name is held by 100 people in the United States SSA records, and 98 of those people were identified as female, then we would infer a 98% probability that the individual in our sample is female. Our approach employs a majority threshold in assigning gender. However, our results are robust to using stricter thresholds, e.g., 80%, 90%, 95%, and indeed they grow stronger. In turn, we use this gender information to construct an indicator of appellant gender, as well as a count-based measure of the number of $FemalePanelists$ that were present.
@@ 3.2.2 Audio Data
Appellant Presents Evidence 0.589 0.492
@@
Appellant Interrupts Panelist 0.335 0.472
Appellant Asks Panelist to Speak Louder 0.065 0.246
Appellant Asks Panelist to Repeat 0.348 0.476
Appellant Sounds Confident 0.896 0.305
Appellant Sounds Hostile 0.111 0.314
@@
Appellant Sounds Angry 0.110 0.312
@@
Panelist Disagrees with Appellant 0.327 0.469
Panelist Raises Voice 0.000 0.016
Panelist Asks Questions 0.845 0.361
Panelist Interrupts Appellant 0.104 0.305
Panelist Asks Appellant to Speak Louder 0.019 0.136
Panelist Asks Appellant to Repeat 0.133 0.339
Panelist Sounds Confident 0.999 0.024
Panelist Sounds Hostile 0.000 0.015
@@ 4 Descriptive Evidence
Assuming the assignment of ARB panelists is a true experiment, we can perform a simple OLS regression, relating the AnySaved dummy to a binary indicator of whether the ARB panel chair is female, absent any controls, and the result will bear a plausibly causal interpretation. Doing so, we obtain the results in Table 4, where we observe evidence of a negative gender concordance effect for female appellants, yet no evidence for any such effect among male appellants. [^13] We have also evaluated these estimates employing randomization inference (Young, 2019), to ensure robustness. Permuting the Female Panel Chair variable 10,000 times across our entire estimation sample and repeating our pair of estimations (one for male appellants and another for female appellants) with each pass, we obtain empirical p-values that indicate a highly significant effect in the case of female appellants (p ¡ 0.001), yet statistically insignificant effects in the case of male appellants (p ¿ 0.10). Further, we obtain an empirical p-value for the difference between the two estimates that also indicates statistical significance (p = 0.0238).
@@
We depict the resulting estimates in Figure 1, ordering panelists along the horizontal axis from least to most likely to adjudicate savings. The average probability that savings are adjudicated spans a range of approximately 40 percentage points, from the most ‘critical’ to the most ‘favorable’ ARB panelist.
@@
Even if ARB panelists exhibited no systematic bias in their decisions, a normal distribution of savings adjudication rates across panelists might nonetheless be expected due to random variation in case characteristics. Accordingly, to explore whether the observed distribution truly reflects panelist-specific biases in decision-making (i.e., whether the distribution truly differs from what we might expect due to chance), we perform a permutation test (Good, 2013). We implement our test via a repeated shuffling of the hearing outcome variable (i.e., AnySaved) over hearings to recover artificial, simulated ARB panelist-specific estimates of savings adjudication rates under the null hypothesis that adjudication rates are, in fact, homogeneous across panelists. That is, by breaking the relationship between the outcome and assigned ARB panelist identities, we effectively simulate the null distribution of our panelist-specific test statistics distributing the panelist-specific bias uniformly across across panelists. This approach effectively retains the structure of panelist case assignments while enabling us to test the null hypothesis that panelists have no systematic effect on outcomes.
@@
In Figure 2, we plot the true set of estimates in red, atop the simulated estimates from 1,000 randomly shuffled samples shown in gray. The simulated distributions show what would result from the objective cases – i.e., the distribution of savings if ARB panelists did not exhibit systematic bias. It is readily apparent that the true distribution has fatter tails than any simulated distribution, such that the dispersion in panelists’ rates of savings adjudication is wider than can be explained by chance alone.
Figure 2: Permutation test evaluating panelist bias.
@@ 5 Main Regression Results
That said, it is important to also recognize that several administrative restrictions limit how property taxes can be adjusted. These restrictions influence the incentives for households to appeal an appraisal, as well as the potential savings households may obtain from an appeal. Further, these restrictions make it difficult to evaluate the temporal evolution of appeals and savings arising from them (Giaccobasso et al., 2025). Most notably, property tax increases are capped at 10% annually. Households already under the cap have less incentive to appeal since their efforts would not immediately reduce their property taxes.
We next explore our measure of the dollar amount saved. As noted earlier, these estimations employ the logarithm of the amount saved as the outcome. The coefficient estimates are thus interpretable as approximate percentage effects. Considering Table 6, we observe the same general pattern of effects as in Table 5. Whereas female homeowners’ case outcomes differ systematically depending on the gender of the ARB panelists, outcomes for male homeowners do not. Considering our estimations in column 1, which employs the same right-hand side model as in column 1 of Table 5, we see that female homeowners save approximately 12.2% less (more) when faced with a female (male) ARB panel chair.151515exp(-0.13) - 1 = -0.122 Further, we see that the amount saved declines with the addition of more female panelists, such that the difference in home value reduction obtained by female homeowners when facing an all-female versus all-male panel is approximately 32.8%.161616exp(-0.397) - 1 = -0.328 Incorporating account fixed effects, we observe consistent findings in the case of ARB chair gender, though statistical significance declines slightly.
@@ 6 Audio Analysis
We next leverage a multi-modal large language model (LLM), specifically Gemini 1.5 Flash, to process all audio recordings associated with nearly all owner-attended ARB hearings in our original sample. Ultimately, we work with a subset of 80,197 audio files, namely those we can reliably match to hearings from our original sample of structured hearing administrative records. We are unable to arrive at matches for all audio files and all structured hearing records because some structured hearing records in our sample do not have associated audio recordings. Further, in some cases, we have multiple structured hearing records for the same property in the same tax year (e.g., if separate hearings were held to adjudicate different protest reasons). In such cases, we lack any means of correctly discerning which audio file is associated with which structured hearing record. In all such cases, we omit the relevant audio files and structured hearing records from our analyses of in-hearing behavior.
Employing Gemini, we construct measures that quantify appellant behavior during a hearing. Using this data, we hope to understand i) any systematic differences that female homeowners may exhibit depending on the gender composition of the panel they face, and ii) any systematic differences in the relationship between homeowners’ behavior and hearing outcomes, depending on ARB panelist gender.
As noted earlier, we consider a measure of hearing length (in seconds), extracted from the metadata of each audio recording. Further, we consider Gemini’s assessment of the fraction of time an appellant spent talking during a hearing, binary indicators of how the appellant sounds, e.g., confident, frustrated, nervous, defensive, etc., and binary indicators of appellant actions during the hearing, e.g., referencing evidence, raising their voice, interrupting a panelist, disagreeing with a panelist, asking questions, asking a panelist to speak louder, or asking a panelist to repeat themselves.